If n=13, x-bar=48, and s=9, find the margin of error at a 80% confidence level

Give your answer to two decimal places.
Answer #1

Solution :

Given that,

t /2,df = 1.356

Margin of error = E = t/2,df * (s /n)

= 1.356 * (9 / 13)

Margin of error = E = 3.38